SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Summary of Prototype II Running

Extracted from proto2data.html 23 July 2001
Information last updated 25 Jun 1998 by Masahiro Morii <masahiro@slac.stanford.edu>

Phase 1
Colorado
Phase 2
SLAC + single-hit TDC
Phase 3
SLAC + multi-hit TDC
Phase 4
FEA
Phase 5
FEA + new DCAC

See also Proto-II Data Structure, Proto-II Cosmic Run Procedure and Proto-II TDC Threshold.


Phase 1 - Runs 381-394

Phase 1 data were recorded using the Colorado preamplifier boards. The discriminator threshold was fixed at 300mV.
Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
381  31-Jan-97/11:35      98  QX0142.1    ---   Calibration
382  31-Jan-97/13:30     870  QX0142.2   1919V  Cosmic, HV -5%
383  31-Jan-97/14:11     430  QX0142.3   1919V  (trigger problem?)
384  31-Jan-97/14:37      57  QX0142.4   1919V  (trigger problem?)
385  31-Jan-97/15:04    1490  QX0142.5   1919V
386  31-Jan-97/17:23   25051  QX0142.6   1919V
387  01-Feb-97/11:16    9772  QX0142.7   1919V
388  01-Feb-97/18:07      16  QX0142.8   1919V  (trigger problem?)
389  01-Feb-97/18:09   24171  QX0142.9   1919V
390  02-Feb-97/11:01    8739  QX0142.10  1919V
391  02-Feb-97/17:16   22653  QX0142.11  1919V
392  03-Feb-97/19:53   19167  QX0142.12  2020V  Cosmic, HV "nominal"
393  04-Feb-97/19:34   22994  QX0142.13  2020V
394  05-Feb-97/19:05   20820  QX0142.14  2020V
===========================================================================
Runs 381 onward were recorded after "fixing" the TDCs by distributing common stop signal at 950ns after each trigger. All TDCs should be working. Event flags are incorrect; ignore them.

Run 381 is a calibration run (but flagged as cosmic data in the file).

Runs with smaller sizes were stopped because of suspected trigger problems (e.g. trigger rates were too high). No clear problems were found, but they may well be discarded.

HV setting for runs 382-391 (called "-5%") was: 1919V on anodes, 865V and 333V on guard wires. This voltage, 5% lower than the nominal voltage calculated using DCSIM, was found to be close to the nominal gas gain of 5x10^4.

HV setting for runs 392-394 (called "nominal") was 2020V on anodes, 910V and 350V on guard. This is the nominal voltage calculated by DCSIM. The gas gain was measured to be about 1-1.2x10^5.

Phase 2 - Runs 682-706

Phase 2 data were recorded using
  • New preamplifier board. Colorado -> SLAC.
    This means different threshold and different t0.
  • DAQ software mods. Calibration and slow data integrated.
Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
682  26-Mar-97/19:43   20003  QX0142.15  2020V
684  27-Mar-97/18:21   20187  QX0142.16  2020V  slow data included
685  28-Mar-97/18:59   21500  QX0142.17  2020V
686  29-Mar-97/10:38   42064  QX0142.18  2020V
689  30-Mar-97/17:54   21965  QX0142.19  2020V
690  31-Mar-97/19:20   20479  QX0142.20  2020V
691  01-Apr-97/18:23   21429  QX0142.21  2020V
700  02-Apr-97/18:30   21267  QX0142.22  2020V  HV -> slow data
702  03-Apr-97/18:31   20657  QX0142.23  2020V
703  04-Apr-97/18:19   24747  QX0142.24  2020V
705  05-Apr-97/12:14   33398  QX0142.25  2020V
706  06-Apr-97/13:36   27418  QX0142.26  2020V
===========================================================================
HV setting is now 2020V/910V/350V by default. The gas gain is re-estimated to be 1.8x10^5 on average, although there is significant difference between layers. Discriminator threshold was set to 260mV. This corresponds to about 1.5 electrons at the above HV setting.

Calibration scheme has changed so that only 1 out of 4 channels per preamp chip is calibrated in each event. This makes 4 (channels) x 10 (delay) x 10 = 400 events for the calibration at the beginning of each run.

Slow control data (temperature, humidity, etc.) have been incorporated into the data stream starting from run 684. The SlowData structure changed (unfortunately):

  • Runs 684-691: 16 variables. There were three variables for HV, which were not read and filled with zeros.
  • Runs 700-706: 17 variables. There were four variables for HV, which were actually filled. The additional one (cathode 4) is for the outer guard (910V) wires.
The latter structure will (hopefully) be kept during the next data taking period.

Data files for runs 689 onward are smaller by a factor 1/3 because channels with no TDC hits were suppressed. (We should've done this long ago...)

Data taking was stopped after run 706 in order to move the chamber from B288 to B273. Total number of triggers = 295,114.

Phase 3 - Runs 847-959

Phase 3 data were recorded using the SLAC preamplifier board and LeCroy multi-hit TDCs. See Changes between run 706 and run847.
Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
847  30-May-97/12:05    3772  QX0142.27  2020V  short waveform
849  30-May-97/19:40   12422  QX0142.28  2020V
850  31-May-97/16:02   14105  QX0142.29  2020V  FADC problem
851  01-Jun-97/15:02   11424  QX0142.30  2020V  FADC problem
852  02-Jun-97/09:22   15283  QX0142.31  2020V  FADC problem
853  03-Jun-97/10:04   15311  QX0142.32  2020V  FADC problem
865  04-Jun-97/20:02    9356  QX0142.33  2020V  FADC problem
899  05-Jun-97/18:47    9789  QX0142.34  2020V
920  07-Jun-97/21:55    8623  QX0142.35  2020V
921  08-Jun-97/11:31   19207  QX0142.36  2020V
922  09-Jun-97/18:26   23751  QX0142.37  2020V  unusually long run
924  11-Jun-97/13:40    9695  QX0142.38  2020V  Trig counters moved backward
925  12-Jun-97/18:45   15677  QX0142.39  2020V
926  14-Jun-97/19:46   14814  QX0142.40  2020V
927  16-Jun-97/18:11    6389  QX0142.41  2020V
928  17-Jun-97/14:45    7451  QX0142.42  2020V  
930  18-Jun-97/15:48   16013  QX0142.43  2020V  lower PMT drawing large current
931  23-Jun-97/09:30    7269  QX0142.44  2020V
932  25-Jun-97/14:15   14761  QX0142.45  2020V  Trig counters moved to middle
935  27-Jun-97/10:50    1403  QX0142.46  2020V  
949  27-Jun-97/17:15   20628  QX0142.47  2020V  pedestal problem
952  30-Jun-97/09:23    1618  QX0142.49  2020V  
959  30-Jun-97/18:05    5485  QX0142.48  2020V
===========================================================================
HV (2020V/910V/350V) and discriminator threshold (260mV) remain the same as the previous period. Calibration scheme is also kept the same, although the multi-hit TDC used in this period has different range and sensitivity from the single-hit TDCs of the previous period.

FADC waveforms were recorded in this period. The nominal length of the recorded waveform is 200ns (10 x 20ns) before and 1000ns (50 x 20ns) after each `signal'. There was a bug in Run 847 which made the waveform shorter, 200ns before and 800ns after each signal.

Runs 850-865 had one FADC module with pedestal problem. The module covers four wires with layer/wire = 14/34, 14/36, 16/34 and 16/36. The waveform from these channels were junk.

After run 922 the iron stack was moved 23 inches toward the electronics. The electronics for layers 5-12 (superlayers 2 and 3) shifted by two cells.

Run 931 contains the last data taken with the iron stack at the end of the prototype. After this run, the stack was moved back to the centre of the chamber and the electronics was appropriately shifted.

Phase 4 - Runs 1175-1182

Phase 4 data were recorded using the preproduction Front-End Assemblies. See Proto-II Cosmic Run Phase 4 for more details.
Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1175 29-Aug-97/15:36     523  QX0142.50  2020V  Vth=300mV, TDC trig data
1176 29-Aug-97/16:51     393  QX0142.51  2020V  Vth=300mV, FADC=3 trig data
1179 29-Aug-97/18:28    4069  QX0142.52  2020V  Vth=300mV
1180 30-Aug-97/11:50    5475  QX0142.53  2020V  Vth=260mV
1181 31-Aug-97/09:42    8343  QX0142.54  2020V  Vth=260mV
1182 01-Sep-97/18:41    3781  QX0142.55  2020V  Vth=260mV
===========================================================================
There are 320 calibration events at the beginning of each run. The sequence of these events are:
For 4 channel mask patterns (0001, 0010, 0100, 1000) {
For 4 strobe timings (17, 34, 51, 68ns) {
Generate 10 pulses
}
For 4 amplitudes (100, 200, 300, 400mV) {
Generate 10 pulses
}
}

Run 1175 and 1176 are short runs for testing the trigger data stream. The discriminator threshold was set to 300mV. The trigger was generated using TDC hits (run 1175) or FADC delta > 3 (run 1176).

Run 1179-1182 are cosmic data for normal resolution studies etc. The discriminator threshold was set to 300mV for run 1179, and to 260mV for run 1180-1182. The electronics was reasonably quiet (only a few noisy channels at 260mV), which was achieved by a dirty workaround using copper tapes et al.

Phase 5 - Runs 1236-1355

Phase 5 data were recorded using the preproduction DCACs and modified FEAs. See Proto-II Cosmic Run Phase 5 for more details.
Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1236 14-Nov-97/19:40    9061  QX0142.56  2020V  Vth=220mV
1237 15-Nov-97/14:37   20510  QX0142.57  2020V  Events after 9898 empty
1238                          QX0142.58  2020V  No reconstructed tracks
1239                          QX0142.59  2020V  DGAIN high, FEA cables swapped
1260                          QX0142.60  2020V  DGAIN low, HV trip: isobutane empty
1302                          QX0142.61  2020V  DGAIN low, BUG: outer elefants killed
1328 25-Nov-97/17:00   11405  QX0142.62  2020V  DGAIN high
1330 26-Nov-97/10:02   17428  QX0142.63  2020V  DGAIN low
1332 28-Nov-97/13:53   14805  QX0142.64  1960V  DGAIN low, gas gain 1/2
===========================================================================

Earlier part of this phase was not very successful due to numerous hardware/software glitches. You should look at the later part, starting from run 1338, for useful data.

Run 1236 and 1237 were taken with two ADBs, one in each FEA. The trigger counters were near the rear end plate to compensate for the limited acceptance of the electronics. The preamplifiers (preproduction) were used with the default gain and full digital bandwidth. The 220mV threshold resulted in a noise hit rate of about 0.3 hit/wire. The data has only 32 calibration events at the beginning of run. This was done similar to the 320 events calibration described above, but with only 1 pulse at each point.

There was a problem during run 1237 after event 9898, which made all subsequent events empty. The run has therefore only 9898 useful events.

The trigger counters were moved back to the center of the chamber before run 1238. This run seems to have had hardware problem; there is no tracks.

Run 1239 was taken with the digital gain raised by a factor of 2. There was a cabling problem which severely limited the geometrical acceptance.

Run 1260 had a gas problem: we ran out of isobutane. The gas gain went high until the HV tripped.

Run 1302 had a problem with software which disabled the outer ADBs. This resulted in almost no geometrical coverage.

Run 1332 was taken with lower HV: anode 1960V, zipper 340 and guard 880V. This should lower the gas gain by a factor of 2.

There was a significant hardware modification between run 1332 and run 1338. The Elefants were found to be measuring the trailing edge of the TDC inputs due to misdefinition of the signal polarity. The signals were flipped so that the leading edge timing would be measured. This problem was affecting all the data since Phase 4.

Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1338 01-Dec-97/21:58    9757  QX0142.65  2020V  DGAIN low
1347 02-Dec-97/16:13   11969  QX0142.66  2020V  DGAIN high
1348 03-Dec-97/11:59    7322  QX0142.67  1960V  DGAIN high
1349 03-Dec-97/23:00    9349  QX0142.68  1960V  DGAIN high
1350 04-Dec-97/19:03    9688  QX0142.69  1900V  DGAIN high, FADC gain=20
1351 05-Dec-97/17:35   11633  QX0142.70  1900V  DGAIN high, FADC gain=20, CALDAC high
1353 06-Dec-97/16:32   12206  QX0142.71  2020V  DGAIN high, FADC gain=20
1354 07-Dec-97/19:06   10148  QX0142.72  1900V  DGAIN high, FADC gain=20
1355 23-Dec-97/14:39    3325  QX0142.73  1960V  trigger recorded
1374 10-Feb-98/18:04   10784  QX0142.74  1960V  New and old ADBs
===========================================================================

Runs 1338 onwards, after the hardware fix, turned out to be very good. Two FEAs (both fully populated with preproduction chips) were used to cover rather large geometrical acceptance. The trigger counter was in the middle. DGAIN of the preamp was switched between low and high, the latter being the expected working point for the real chamber. Three HV points were tried: 2020V as a reference (same as many old runs), 1960V and 1900V for more realistic working points. The absolute gas gain is yet to be determined, but the gain ratios are believed to be about G(1960V)/G(2020)V = 0.60 and G(1900V)/G(2020)V = 0.35.

For most runs, the analog gain of the Elefant was set to 5 (nominal). We tried to change it to 10 for runs 1350 and 1351, where the low HV setting would have make it difficult to measure the waveform. Due to a software problem, we ended up changing the analog gain to 20. The FADC was always set to linear mode (V_top = 4.0V, V_mid = 2.5V, V_bot = 1.0V).

In run 1351, CALDAC was set to the maximum value (4.095V). This was found to reduce the noise by ~20mV for some yet unknown reason. With DGAIN set to high, we could run at threshold = 260mV compared with 300mV in run 1350. It is not yet clear if this really improved the S/N ratio; it might just have reduced the digital gain.

Run 1353 was meant to be a repeat of run 1347, but I forgot to set the FADC gain back to 5. It should not matter as far as the spatial resolution is concerned.

Run 1355 is a short run for testing trigger bits. The condition was: HV=1960V, dgain high, FADC gain (still) 20, threshold 300mV. CALDAC was normal (100mV).

Run 1374 is a test run for the new (3rd generation) ADBs. Two FEA boxes equipped with old (lefti box) and new (right box) ADBs were used to take the data at HV = 1960V and DGAIN = low. The digital gain was not completely uniform: One of the old ADBs had a diode on the DGAIN line making it open when the DGAIN is set low; other ADBs (1 old and 2 new) had 33k resistors to the CMOS driver making the gain a little lower. The threshold was 200mV, CALDAC 4V, FADC gain 10 (I believe).

Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1390 28-Feb-98/17:38   11160  QX0142.75  1960V  \
1391 01-Mar-98/11:43   13433  QX0143.1   2020V   \
1392 02-Mar-98/17:48   13098  QX0143.2   1900V    \ Junk
1393 03-Mar-98/18:53   13703  QX0143.3   1960V    / Threshold set to 500mV
1394 04-Mar-98/21:42   15971  QX0143.4   2020V   /
1395 05-Mar-98/18:45   12730  QX0143.5   1900V  /
1397 06-Mar-98/19:23   15309  QX0143.6   1960V  \  3rd generation FEA
1398 07-Mar-98/17:58   13565  QX0143.7   2020V   > Pedestal calibration
1399 08-Mar-98/14:15   13582  QX0143.8   1900V  /  Bilinear FADC
1410 04-Apr-98/12:58   18870  QX0143.9   1960V  FADC gain 10
1411 05-Apr-98/14:49   14414  QX0143.10  1960V  FADC gain 5
===========================================================================

Runs 1390 onwards used two pre-production FEAs fully loaded with 3rd generation ADBs. The running conditions were chosen to be as close as possible to what we will use on the real chamber. DGAIN = low. threshold = 200mV. FADC was set to bi-linear mode, i.e., Vtop = 4.0V, Vmid = 1.75V, Vbot = 1.0V. Vref was set to 1.1V to bring the pedestal around 5 FADC count.

In order to facilitate the determination of the FADC pedestal in each event, the timing of the Level-1 Accept was delayed by 16 sysclks (=267ns) compared to the previous runs. This should move the waveforms by 4 FADC samples, leaving more room before the signal. A side effect of this change is that the measured hit time is now shifted by 267ns and the T0 must be retuned.

We also started to do pedestal calibration at the beginning of each run. This consists of 110 events with no signal (readout sparsification was disabled, of course). During the pedestal calibration, Vref was scanned within ±50mV of the nominal value in 11 steps, changing the pedestal value by ±2-3 counts. Averaging the pedestal values over these 11 Vref values should wash out the quantization error and give us a better measurement of the true pedestal.

Unfortunately, a bug was introduced in the start-of-run procedure which set the threshold to 500mV for the first 6 runs of this period. Runs 1390-1395 are therefore junk. The bug was fixed before Run 1397.

From Run 1397 onwards, the TDC calibration was done with the "small-signal" calibration mode, i.e. SEL33X = 1, and CALDAC set to 100mV. The FADC calibration was done in the "large-signal" mode with CALDAC scanning from 50mV to 200mV in 50mV steps.

Data taking resumed from Run 1410 after a month of break. The condition was kept same: 3rd generation ADB with preproduction DCACs, "low" DGAIN, threshold at 200mV. HV was kept at the nominal 1960V.

Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1420 11-Apr-98/22:41   10676  QX0143.11  2020V  \ New DCAC, thresh = 200mV
1421 12-Apr-98/21:57   10803  QX0143.12  2020V  / Junk runs
1431 13-Apr-98/20:28   11584  QX0143.13  2020V  thresh 200mV, gas problem?
1433 14-Apr-98/18:46   11436  QX0143.14  1960V  thresh 180mV, 2 final ADBs
1435 16-Apr-98/00:18   10506  QX0143.15  1960V  thresh 180mV, 3 final ADBs
1437 17-Apr-98/19:10   24295  QX0143.16  2020V  \  4 final ADBs
1438 18-Apr-98/16:36   13624  QX0143.17  1960V   > thresh 180mV
1439 19-Apr-98/11:14   20221  QX0143.18  1960V  /
===========================================================================

Runs 1420 onwards used Production DCAC. First two runs, 1420 and 1421, were taken before the chip's peculiarities were understood, and are therefore junk. HV was set to 2020V and the threshold was 200mV.

Significant progress had been made in understanding the final DCAC before Run 1431. The stability of the chips had been greatly improved by setting the common-mode voltage of the threshold inputs at 2.75V instead of 2.4V.

The first run after this breakthrough, Run 1431, turned out to be a disappointement. We suspect that the gas mixture was not right at the beginning (gas flow was set very low), and the calibration changed during the run. There was also a problem with one ADB in which the FET bias voltage was wrong and the FADC waveforms were lost. HV was set to 2020V and the threshold was 200mV.

Runs 1433 and 1435 were taken at HV = 1960V and the threshold of 180mV. The data quality turned out to be good. Up to this point, we had a mixture of final and pre-final ADBs. The threshold was limited by the pre-final ADBs which showed somewhat poorer signal-to-noise ratios.

Run 1437 onwards were taken with all final ADBs. At this point, all the components of the FEAs were final. The digital-gain control resistors (33kohm on DGAIN pin of DCAC) were removed, resulting in about 20% increase in the digital gain. The threshold remained at 180mV. HV was set to 2020V for Run 1437, 1960V for Runs 1438 and 1439.

Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1510 03-Jun-98/11:16   27508  QX0413.19  2020V  He-CO2-C4H10, Ele gain 5
1614 17-Jun-98/19:22   22433  QX0413.20  2020V  He-CO2-C4H10, Ele gain 10
1640 19-Jun-98/17:11   29059  QX0413.21  2100V  He-CO2-C4H10, Ele gain 10
===========================================================================
Run 1510 and 1614 were taken with a non-flammable gas mixture, He-CO2-C4H10. HV was set to 2020V, discriminator threshold 180mV. Elefant analog gain was 5 for Run 1510 and 10 for Run 1614. Run 1640 was taken with HV=2100V.
Run  Start date/time  Events  Tape.seq   HV     Comments
===========================================================================
1792 05-Aug-98/09:07     759             1960V  Gas test run
===========================================================================
Run 1792 was taken for checking the quality of a new isobutane bottle. The ``LAB'' DAQ system with one ROM was used to record this data. The discriminator threshold was 250mV to suppress the excess noise seen only with ROM+DIOM on Proto-II. The gas mixture is close to 80:20, but may not be very accurate. The FADC gain was set to 10.